OpenStudy (anonymous):

What is the solution of ?Answer {-4, 5} {-2} {-2, 3} {3}

OpenStudy (anonymous):



OpenStudy (anonymous):

equal 5 @jennychan12

OpenStudy (jennychan12):

aishhh absolute value 2x-1 = 5 -2x+1 = 5 solve for x in both

OpenStudy (anonymous):

OK, so the question is: |2x-1| =5 I'd break that down: 2x -1 =5; or 2x - 1 = -5 Let's do the first part first. 2x-1=5 2x=6 x=3 Now let's do the second part: 2x-1=-5 2x=-4 x=-2 So the final answer is x = 3 or -2 (You can also substitute each one in the original equation to check.) Hope that helps!

OpenStudy (anonymous):

So -2 and 3 ?

OpenStudy (jennychan12):

yepp

OpenStudy (anonymous):

Yes, it would be the third one of the choices shown.